• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Sirca

    Log entry time 12:07:34 on October25,2001

    Entry number 74971

    keyword=Right Q1 cryo problem fixed

    Hall access started at 09:30 to reset two fire alarms and to check the cryogenics
    of Right Q1. Reason for low cryo levels was a closed valve. Reopened it, Mark
    will monitor helium levels for the next few hours, but it seems fine OK.
    Tested Right Q1 with 1400 A and set it to momentum 1.066 GeV/c afterwards.
    Hall access finished at 10:30.